Solve for x
-8x+14=-2(4x-7)
: Solve for x
-8x+14=-2(4x-7)

Answer by jim_thompson5910(9392)


-8x+14=-2(4x-7) Start with the given equation.


-8x+14=-8x+14 Distribute.


-8x=-8x+14-14 Subtract 14 from both sides.


-8x+8x=14-14 Add 8x to both sides.


0x=14-14 Combine like terms on the left side.


0x=0 Combine like terms on the right side.


0=0 Simplify.


Since this equation is always true for any x value, this means x can equal any number. So there are an infinite number of solutions.